OpenStudy (anonymous):

Derive the trigonometric addition formula for sine :

OpenStudy (anonymous):

\[\sin(a+b)=sinacosb+cosasinb\]

OpenStudy (anonymous):

also, derive these identities using the addition or subtraction formulas for sine or cosine : \[sinasinb=\frac{ 1 }{ 2 }(\cos(a-b)-\cos(a+b))\]

OpenStudy (irishboy123):

easiest i know is: \( e^{i \ a} = (cos a + i \ sin a)\) \( e^{i \ b} = (cosb + i \ sin b)\) \( e^{i \ a} . e^{i \ b} = e^{i \ (a+b)} = \) then multiply out and equate the real and complex bits for the second bit, just churn the formula. eg expand the RHS using the cos (a+b) and cos (a-b).

OpenStudy (irishboy123):

if you post some stuff, i'm sure you will get all the help you need :p

OpenStudy (anonymous):

I've never seen that first formula before . I'm even more confused

OpenStudy (freckles):

you can derive cos(a-b)=cos(a)cos(b)+sin(a)sin(b) using the distance formula then you can use cos(a-b)=cos(a)cos(b)+sin(a)sin(b) along with cofunction identities to show sin(a+b)=sin(a)cos(b)+sin(b)cos(a)

OpenStudy (anonymous):

what are cofunction identities ?

OpenStudy (anonymous):

This is an algebra 2 class and they threw in a trig section

OpenStudy (freckles):

you are asking to derive the formula sin(a+b)=sin(a)cos(b)+sin(b)cos(a) right?

OpenStudy (anonymous):

yes

OpenStudy (freckles):

hmmm maybe you don't remember the name of the identity but you remember that cos(90-A)=sin(A) and sin(90-A=cos(A)? If not you easily show these by drawing right triangles and comparing the above ratios.

OpenStudy (anonymous):

the notes didn't name the identities and I'm sure I learned that . Trigonometry is just really hard for me to understand

OpenStudy (irishboy123):

hi @jxaf sorry for the confusion, i should have asked first what you knew. do you think you could follow this proof, with some guidance. https://gyazo.com/be936132eb2a67f24dbd6841f0cf94a3

OpenStudy (anonymous):

Um, that looks really complex. Do you want a picture of my notes to see what I learned ?

OpenStudy (irishboy123):

yeah, go for it!

OpenStudy (anonymous):

okay it might take a bit cause my computer is really slow

OpenStudy (irishboy123):

take your time.

OpenStudy (irishboy123):

well i've drawn the shape for the proof anyways just in case it comes in useful. i think it's more straightforward that it might appear.